### v1

# Pseudo-real quantum fields

Abstract
We introduce the concept of pseudo-reality for complex numbers. We show that this concept, applied to quantum fields, provides a unifying framework for two distinct approaches to pseudo-Hermitian quantum field theories. The first approach stems from analytically continuing Hermitian theories into the complex plane, while the second is based on constructing them from first principles. The pseudo-reality condition for bosonic fields resolves a long-standing problem with the formulation of gauge theories involving pseudo-Hermitian currents, sheds new light on the resolution of the so-called Hermiticity Puzzle, and leads to a consistent minimal coupling of pseudo-Hermitian quantum field theories to gravity. These results contribute significantly to the ongoing development of the first-principles construction of pseudo-Hermitian quantum field theories, including their formulation in curved spacetimes.
1 Introduction
The success of non-Hermitian quantum mechanics has led to increasing interest in extending these ideas to non-Hermitian quantum field theory, with much of the progress so far focused on $PT$-symmetric quantum field theories, where the Hamiltonian is symmetric under the combined action of parity $P$ and time-reversal $T$. When the $PT$ symmetry is exact, the Hamiltonian has a real spectrum and the theory follows a (pseudo-)unitary evolution, which ensures its physical viability [1, 2, 3, 4, 5, 6] . Well-studied examples include scalar field theories with an imaginary cubic $i\phi^{3}$ [7, 8, 9, 10] or “wrong-sign” $-\phi^{4}$ quartic self-interaction [11, 12, 13].
However, non-Hermitian quantum field theories bring new challenges that are not present in non-Hermitian quantum mechanics. One such issue is the Hermiticity Puzzle [14], which arises when a Hermitian theory with real fields is made non-Hermitian by introducing complex couplings. This modification leads to equations of motion being inconsistent under complex conjugation. Previous resolutions of the Hermiticity Puzzle required real fields to transform non-trivially under complex conjugation [14].
Inconsistent Euler-Lagrange equations are a common feature of non-Hermitian quantum field theories, as first noted in Ref. [15], and are not limited to theories with real fields. These theories also lack Poincaré invariance, and it remains unclear how to construct consistent gauge theories or couple them to gravity. In our previous work [16] (see also Refs. [17, 18, 19]), we traced these problems to a fundamental mismatch: the field $\phi$ and its Hermitian conjugate $\phi^{\dagger}$ transform under different representations of the proper Poincaré group. This mismatch not only breaks Poincaré invariance but also causes $\phi$ and $\phi^{\dagger}$ to evolve under different Hamiltonians, $H$ and $H^{\dagger}$, respectively. To address this, we proposed replacing the Hermitian conjugate field $\phi^{\dagger}$ with a dual field $\tilde{\phi}^{\dagger}$, which transforms under the same representation of the proper Poincaré group as $\phi$ and thus, evolves with the same Hamiltonian.
In this paper, we extend this idea to resolve the Hermiticity Puzzle by introducing the concept of pseudo-reality. Fields that are real in Hermitian theories become pseudo-real in pseudo-Hermitian theories, i.e., rather than being equal to their Hermitian conjugates $\phi=\phi^{\dagger}\in\mathbb{R}$, they are equal to their duals $\phi=\tilde{\phi}^{\dagger}\in\mathbb{C}$. Thus, the fields maintain the same number of degrees of freedom as a real scalar field, while being complex. This approach simultaneously resolves long-standing questions about how to gauge non-Hermitian quantum field theories and couple them to gravity in a way that is consistent with the local symmetries of the Lagrangian and current conservation. Additionally, it unifies two approaches to constructing non-Hermitian quantum field theories found in the literature: one based on the analytic continuation of Hermitian theories into the complex space, and the other based on building the non-Hermitian theory from first principles.
The paper is organised as follows. In Section 2, we introduce the concept of pseudo-reality for complex functions and establish the conditions that a function must satisfy to be pseudo-real, providing examples for illustration. In Section 3, we apply the pseudo-real formalism to the non-Hermitian $i\phi^{3}$ quantum field theory. We demonstrate that not only must the field be complex, but the parity and time-reversal transformations must also be non-trivial. We find the transformation under which the Hamiltonian is pseudo-Hermitian. In Section 4, we extend this approach to higher-spin fields. We show how to consistently gauge a two complex scalar model with non-Hermitian mass mixing and outline how to couple this model to gravity.
2 Definition of pseudo-reality for complex numbers
The number $z$ is said to be real if it satisfies the following condition:
$z={\bar{z}}\qquad\ {\text{[for a real $z$]}}\,.$ (2.1)
The standard definition of reality (2.1) for a complex number $z$ involves only the operation of complex conjugation $z\to{\bar{z}}$. The notion of reality can be generalized if one considers the reality condition with respect to a complex function $f=f(z)$. To this end, we consider a complex map $f:\,\mathbb{C}\mapsto\mathbb{C}$. If the function $f=f(z)$ is analytic and satisfies the property
$f^{(-1)}(z)={\bar{f}}(z)\,,$ (2.2)
valid for a continuous segment of the arguments $z\in\mathbb{C}$, where $\bar{f}$ is the complex conjugate of the map $f$, then $f$ can be used to define a curve in the complex plane via
$z=f({\bar{z}})\,.$ (2.3)
We then call the number $z$ a pseudo-real number with respect to the complex map $f$ ($z\in\mathbb{R}_{f}$), provided the pseudo-reality condition (2.3) is fulfilled and the function $f$ satisfies the consistency condition (2.2). Notice that Eq. (2.3) implies automatically that
$\bar{z}=f^{(-1)}(z)\;,$ (2.4)
which can also be used as a definition of the pseudo-reality, fully consistent with Eq. (2.3).
Condition (2.2) involves the inverse function $f^{(-1)}$ defined via $f^{(-1)}\Bigl{(}f(z)\Bigr{)}=z$. Therefore, the complex conjugation on the right-hand side of Eq. (2.2) acts only on the function $f$, excluding its argument $z$, i.e.,
${\bar{f}}(z)\equiv\overline{f({\bar{z}})}\,.$ (2.5)
Thus, Eq. (2.2) can be understood in terms of the Laurent series of the function $f$ and its inverse $f^{(-1)}$ as follows:
${\rm if}\quad f(z)=\sum_{n\in{\mathbb{Z}}}C_{n}z^{n}\quad{\rm and}\quad f^{(-1)}(z)=\sum_{n\in{\mathbb{Z}}}C^{(-1)}_{n}z^{n}\quad{\rm then}\quad{\overline{C}}_{n}=C^{(-1)}_{n},\ \ \forall n\in{\mathbb{Z}}\,.$ (2.6)
If one writes $z=x+iy$, then Eq. (2.3) gives us the following constraint on the variables $x$ and $y$:
$x=u(x,y)\,,\qquad y=v(x,y)\,,$ (2.7)
where $f(\bar{z})=u(x,y)+iv(x,y)$. The system of two equations (2.7) provides us, in general, with a solution in terms of a list of isolated points. In order to obtain a line, this set of generally nonlinear equations should be degenerate so that these two equations should be satisfied by a single function $y=y(x)$. It is the constraint (2.2) that makes system (2.7) degenerate.
Let us consider now a few simple examples.
  • Take first a linear function $f(z)=C_{0}+C_{1}z$, where $C_{0}$ and $C_{1}$ are complex coefficients. In order for this function to be appropriate for the definition of the pseudo-reality of the complex number $z$ (2.3), the function $f=f(z)$ must satisfy the consistency condition (2.2). The latter condition gives us the solution for the function $f$ in the following form:
    $f(z)=e^{2i\theta}z\,,\qquad\theta\in{\mathbb{R}}\,,$ (2.8)
    thus implying that $C_{0}=0$ and $C_{1}=e^{2i\theta}$, where $\theta\in[0,2\pi)$ is a real number. Then the pseudo-reality condition (2.3) reads as follows:
    $z=e^{2i\theta}{\bar{z}}\,.$ (2.9)
    The standard reality condition (2.1) corresponds to a pseudo-real number that satisfies (2.9) with $\theta=0$. For a general angle variable $\theta$, the solution of the pseudo-reality condition (2.9) defines a straight line in the complex plane, $z=\rho e^{i\theta}$ with $\rho\in\mathbb{R}$. Alternatively, in terms of $z=x+iy$, it gives the line $y=x\tan\theta$ with the inclination of the angle $\theta$ to the abscissa. While trivial, this gives an intuitive example of what is meant by a pseudo-real number: it is one with only one degree of freedom but a non-trivial complex conjugate.
  • For a non-linear function, let us consider a generic perturbative solution in terms of $f(z)=C_{1}z+C_{2}z^{2}$, with $|C_{2}z|\ll|C_{1}|$. Solving Eq. (2.2) for small arguments, one can demonstrate that the solution is:
    $f(z)=e^{2i\theta}z\pm ige^{3i\theta}z^{2}+O(g^{2}z^{3})\,,\qquad g\in{\mathbb{R}}\,,\quad|gz|\ll 1\,.$ (2.10)
    Functions of this form will arise in the case of quantum field theories with non-Hermitian interaction terms, as we will see in the next section in the context of the $i\phi^{3}$ theory.
  • Consider a generic Möbius transformation:
    $f(z)=\frac{az+b}{cz+d}\,,\qquad{\rm with\ the\ inverse}\qquad f^{-1}(z)=\frac{dz-b}{-cz+a}\,.$ (2.11)
    The condition (2.2), which defines the allowed function $f$, gives us the following set of degenerate constraints:
    ${\overline{\mathsf{M}}}=({\rm det}\,{\mathsf{M}}){\mathsf{M}}^{-1}\,,\qquad{\mathsf{M}}=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\,,$ (2.12)
    or $a=\bar{d}$, $b=-\bar{b}$, and $c=-\bar{c}$. The constraint is resolved in the form:
    ${\mathsf{M}}=\begin{pmatrix}a&i\kappa_{1}\\ i\kappa_{2}&a\end{pmatrix}\,,\qquad\kappa_{1,2}\in{\mathbb{R}}\,,\quad a\in{\mathbb{C}}\,.$ (2.13)
    Thus, the Möbius function (2.11) with the coefficients (2.13) satisfies the requirement (2.2) imposed on the function $f=f(z)$ to be used in the definition (2.2) of the pseudo-reality for a complex number $z$. Then, the linear manifold identified by Eq. (2.3) has the following explicit form:
    $(x-x_{0})^{2}+(y-y_{0})^{2}=r_{0}^{2}\,.$ (2.14)
    This formula describes a circle centered at the position $(x,y)=(x_{0},y_{0})$ with the radius $r_{0}$, related to the components of the matrix ${\mathsf{M}}$, Eq. (2.13), by the following equations, respectively:
    $x_{0}=\frac{a_{2}}{\kappa_{2}}\,,\qquad y_{0}=-\frac{a_{1}}{\kappa_{2}}\,,\qquad r_{0}^{2}=x_{0}^{2}+y^{2}_{0}-\frac{\kappa_{1}}{\kappa_{2}}\,.$ (2.15)
    Notice that, at the parameter subspace with $r^{2}_{0}<0$, no solution exists.
3 Pseudo-real $i\phi^{3}$ scalar field theory
The $i\phi^{3}$ theory, extensively studied by Bender and collaborators [7, 8, 9, 10], is the quantum field theory counterpart of the quantum mechanical model $H=p^{2}+ix^{3}$, where the spectrum was shown to be entirely real [1, 20]. Unlike the Hermitian $g\phi^{3}$ theory, which is unphysical due to its unbounded energy, the non-Hermitian $ig\phi^{3}$ theory has a bounded energy spectrum that remains real under exact $PT$ symmetry.
In this section, we revisit the $i\phi^{3}$ theory using the first-principles approach introduced in our previous work [16]. We show that extending the coupling $g\mapsto ig$ into the complex plane requires the field $\phi$ to be complex and assuming ‘standard’ parity and time-reversal transformations leads to a contradiction. To resolve this, we formulate the theory in terms of the field $\phi$ and its dual $\tilde{\phi}^{\dagger}$, both evolving with the same Hamiltonian. We introduce the concept of pseudo-reality, where $\phi$ is no longer real but a complex pseudo-real field. We look for a ‘parity-like’ transformation $\eta_{P}$ and a ‘time-reversal-like’ transformation $\eta_{T}$ with respect to which the Hamiltonian is pseudo-Hermitian and $PT$-symmetric, making this a self-consistent formulation.
The Lagrangian of the $i\phi^{3}$ theory is often written as
$\mathcal{L}(x)=\frac{1}{2}\partial_{\mu}\phi(x)\partial^{\mu}\phi(x)-\frac{1}{2}m^{2}\phi^{2}(x)-\frac{ig}{3!}\phi^{3}(x)\;,$ (3.1)
where $m^{2}>0$ and $g\in\mathbb{R}$ is some real coupling.
In the ‘standard’ approach, $\phi$ is considered to be a real scalar field, so that non-Hermiticity arises through the complex coupling $ig$:
$\mathcal{L}^{\dagger}(x)=\frac{1}{2}\partial_{\mu}\phi(x)\partial^{\mu}\phi(x)-\frac{1}{2}m^{2}\phi^{2}(x)+\frac{ig}{3!}\phi^{3}(x)\;.$ (3.2)
The non-Hermitian Lagrangian in Eq. (3.1) is $PT$-symmetric under the ‘standard’ parity and time-reversal transformations, where $\phi$ transforms as a pseudo-scalar field:
$P:$ $\>\>\>\>\phi(x)\longmapsto\phi^{P}(x_{P})=-\phi(x)\>,\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\hat{\mathcal{P}}\hat{\phi}(x_{P})\hat{\mathcal{P}}^{-1}=-\hat{\phi}(x)\>,$ (3.3a)
$T:$ $\>\>\>\>\phi(x)\longmapsto\phi^{T}(x_{T})=\phi(x)\>,\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\hat{\mathcal{T}}\hat{\phi}(x_{T})\hat{\mathcal{T}}^{-1}=\hat{\phi}(x)\>,$ (3.3b)
$PT:$ $\>\>\mathscr{L}(x)\longmapsto\mathscr{L}^{PT}(x_{PT})=\mathscr{L}(x)\>,\>\>\>\>\>\>\>\>\>\>(\hat{\mathcal{P}}\hat{\mathcal{T}})\hat{\mathscr{L}}(x_{PT})(\hat{\mathcal{P}}\hat{\mathcal{T}})^{-1}=\hat{\mathscr{L}}(x)\>.$ (3.3c)
Here, we consider the transformation of the field function $\phi$ on the left and the field operator $\hat{\phi}$ on the right.
The existence of an anti-linear symmetry implies that the Lagrangian is pseudo-Hermitian [6, 21], i.e., there exists some linear transformation $\eta$ such that it takes the Lagrangian to its Hermitian conjugate:
$\eta:\>\mathscr{L}(x)\longmapsto\mathscr{L}^{\eta}(x_{\eta})=\mathscr{L}^{\dagger}(x)\>,\>\>\>\>\>\>\hat{\eta}\hat{\mathscr{L}}(x_{\eta})\hat{\eta}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)\>.$ (3.4)
Here, we take into account that $\eta:x\mapsto x_{\eta}$ could be a coordinate transformation (e.g., parity, $\mathcal{P}:x\mapsto x_{P}$).
Indeed, the Lagrangian in Eq. (3.1) is pseudo-Hermitian with respect to the ‘standard’ parity transformation in Eq. (3.3a):
$P:\>\mathscr{L}(x)\longmapsto\mathscr{L}^{P}(x_{P})=\mathscr{L}^{\dagger}(x)\>,\text{\>\>\>\>\>\>\>}\hat{\mathcal{P}}\hat{\mathscr{L}}(x_{P})\hat{\mathcal{P}}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)\>.$ (3.5)
It is also anti-pseudo-Hermitian with respect to the ‘standard’ time-reversal transformation in Eq. (3.3b):
$T:\>\mathscr{L}(x)\longmapsto\mathscr{L}^{T}(x_{T})=\mathscr{L}^{\dagger}(x)\>,\>\>\>\>\>\>\>\>\>\hat{\mathcal{T}}\hat{\mathscr{L}}(x_{T})\hat{\mathcal{T}}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)\>.$ (3.6)
We refer to the system satisfying Eqs. (3.5) and (3.6) as possessing the property of ‘anti-pseudo-Hermiticity’, since the time-reversal operator $T$ is anti-linear [6]. The corresponding Hamiltonian
$H=\int{\rm d}^{3}\vec{x}\>\quantity(\frac{1}{2}\dot{\phi}^{2}+\frac{1}{2}(\nabla\phi)^{2}+\frac{1}{2}m^{2}\phi^{2}+\frac{ig}{3!}\phi^{3})$ (3.7)
is also parity-pseudo-Hermitian and time-reversal-anti-pseudo-Hermitian:
$P:\>$ $H\longmapsto H^{\dagger}\>,\text{\>\>\>\>\>\>}\>\>\>\hat{\mathcal{P}}\hat{H}\hat{\mathcal{P}}^{-1}=\hat{H}^{\dagger}\>,$ (3.8a)
$T:\>$ $H\longmapsto H^{\dagger}\>,\text{\>\>\>\>\>\>\>}\>\>\hat{\mathcal{T}}\hat{H}\hat{\mathcal{T}}^{-1}=\hat{H}^{\dagger}\>.$ (3.8b)
3.1 Issues with the standard approach
The ‘standard’ approach described above has several issues. First, the field $\phi$ cannot be a real scalar field. This can be seen directly from the equations of motion
$\Box_{x}\phi(x)+m^{2}\phi(x)+\frac{ig}{2}\phi(x)^{2}=0\;,$ (3.9a)
$\Box_{x}\phi^{{\dagger}}(x)+m^{2}\phi^{{\dagger}}(x)-\frac{ig}{2}\phi^{{\dagger}2}(x)=0\;,$ (3.9b)
which are not compatible under Hermitian conjugation if we take $\phi=\phi^{{\dagger}}$. Indeed, if we solve these equations perturbatively in $g$, we see that $\phi$ differs from a real field already at order $g$:
$\phi(x)=\phi_{0}(x)+\frac{ig}{2}\int{\rm d}^{4}y\>G_{0}(x,y)\phi_{0}^{2}(y)+\mathcal{O}(g^{2})\>.$ (3.10)
Here, $\phi_{0}$ is the free real Klein-Gordon scalar field and $G_{0}(x,y)$ is the free Klein-Gordon Green’s function, i.e.,
$\quantity(\Box_{x}+m^{2})\phi_{0}(x)$ $=0\;,$ (3.11)
$(\Box_{x}+m^{2})G_{0}(x,y)$ $=-\delta^{(4)}(x-y)\;.$ (3.12)
We can also look at Hamilton’s equations of motion for the field function and the field operator
$\quantity{\phi(x),H}_{\rm PB}=\partial_{t}\phi(x)\text{\>\>\>and\>\>\>}[\hat{\phi}(x),\hat{H}]=i\partial_{t}\hat{\phi}(x)\>,$ (3.13)
where “PB” indicated the Poisson bracket. By taking their Hermitian conjugates, we see that $\phi^{\dagger}$ does not evolve with the same Hamiltonian as $\phi$
$\quantity{\phi^{\dagger}(x),H^{\dagger}}_{\rm PB}=\partial_{t}\phi^{\dagger}(x)\text{\>\>\>and\>\>\>}[\hat{\phi}^{\dagger}(x),\hat{H}^{\dagger}]=i\partial_{t}\hat{\phi}^{\dagger}(x)\>.$ (3.14)
Hence, for any canonical formalism with a non-Hermitian Hamiltonian, the fields must be complex. This observation also shows that the Hermitian-conjugate field $\phi^{\dagger}$ does not transform under the same representation of time translations as $\phi$, which we discuss in the following subsection in the context of the proper Poincaré group. Moreover, the introduction of a complex scalar field doubles the degrees of freedom of the original theory.
The second issue arises from assuming the ‘standard’ parity and time-reversal transformations in Eqs. (3.3a) and (3.3b). Using pseudo-Hermiticity of the Hamiltonian in Eqs. (3.8a) and (3.8b), we find that Hamilton’s equations transform inconsistently:
$P:\>\quantity{\phi(x),H}_{\rm PB}=\partial_{t}\phi(x)$$\longmapsto\quantity{\phi(x),H^{\dagger}}_{\rm PB}=\partial_{t}\phi(x)\>,$ (3.15a)
$T:\>\quantity{\phi(x),H}_{\rm PB}=\partial_{t}\phi(x)$$\longmapsto\quantity{\phi(x),H^{\dagger}}_{\rm PB}=\partial_{t}\phi(x)\>,$ (3.15b)
$\hat{\mathcal{P}}[\hat{\phi}(x),\hat{H}]\hat{\mathcal{P}}^{-1}=i\hat{\mathcal{P}}\partial_{t}\hat{\phi}(x)\hat{\mathcal{P}}^{-1}$ $\longmapsto[\hat{\phi}(x),\hat{H}^{\dagger}]=i\partial_{t}\hat{\phi}(x)\>,$ (3.15c)
$\hat{\mathcal{T}}[\hat{\phi}(x),\hat{H}]\hat{\mathcal{T}}^{-1}=-i\hat{\mathcal{T}}\partial_{t}\hat{\phi}(x)\hat{\mathcal{T}}^{-1}$ $\longmapsto[\hat{\phi}(x),\hat{H}^{\dagger}]=i\partial_{t}\hat{\phi}(x)\>.$ (3.15d)
Hence, the parity and time-reversal transformations which make the theory pseudo-Hermitian and $PT$-symmetric are not the ‘standard’ transformations but instead are non-trivial (see Refs. [17, 18, 16]). We will find them explicitly in the next subsections.
3.2 Pseudo-Hermitian formalism
For a quantum field theory with a non-Hermitian Hamiltonian, the field $\phi$ and its Hermitian conjugate $\phi^{\dagger}$ transform under different representations of the proper Poincaré group. Thus, not only do they not evolve with the same Hamiltonian, but also the Lagrangian composed of $\phi$ and $\phi^{\dagger}$ is not Poincaré invariant [16]. Instead in Ref. [16] (see also Refs. [17, 18, 19]), we propose to build the Lagrangian from the field $\phi$ and its ‘dual’ field $\tilde{\phi}^{\dagger}$, which transform in the same representation of the proper Poincaré group.
Given that the Lagrangian is pseudo-Hermitian with respect to some Hermitian operator $\hat{\eta}$, i.e., $\hat{\eta}\hat{\mathscr{L}}(x_{\eta})\hat{\eta}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)$ and $\hat{\eta}^{\dagger}=\hat{\eta}$, the ‘dual’ field operator $\hat{\tilde{\phi}}^{\dagger}$ has a general form [16]
$\hat{\tilde{\phi}}^{{\dagger}}(x)=\hat{\eta}^{-1}\hat{\phi}^{\dagger}(x_{\eta})\hat{\eta}\>\pi\>.$ (3.16)
Here, $x_{\eta}$ is the coordinate transformation under $\hat{\eta}$ (e.g., parity $x_{P}$). For an $n$-component field $\hat{\phi}^{a}$, where $a\in\{1,\cdots,n\}$, $\pi$ is a Hermitian $n\times n$ matrix. Hence, for a $1$-component scalar field, $\pi\in\mathbb{R}$ is some real constant to be determined.
The ‘dual’ field operator evolves with the same Hamiltonian as $\phi$ and thus has the same equation of motion:
$[\hat{\tilde{\phi}}^{{\dagger}}(x),\hat{H}]=i\partial_{t}\hat{\tilde{\phi}}^{{\dagger}}(x)\>,\qquad\ \Box_{x}\hat{\tilde{\phi}}^{{\dagger}}(x)+m^{2}\hat{\tilde{\phi}}^{{\dagger}}(x)+\frac{ig}{2}\hat{\tilde{\phi}}^{{\dagger}^{2}}(x)=0\>.$ (3.17)
We can also write the ‘dual’ field as a field function in terms of the map $\eta$ (3.4):
$\tilde{\phi}(x)\coloneqq\pi\eta[\phi(x)]\>\implies\>\tilde{\phi}^{\dagger}(x)=\eta^{\dagger}[\phi^{\dagger}(x)]\pi\>,$ (3.18)
which relates the transformation $\eta$ to the operator $\hat{\eta}$ via:
$\eta[\hat{\phi}(x)]=\hat{\eta}\hat{\phi}(x_{\eta})\hat{\eta}^{-1}\>.$ (3.19)
We can take Eq. (3.19) as a definition of the functional map $\eta$. This map is linear, and the Hermiticity of the operator $\hat{\eta}$ implies that the transformation $\eta^{\dagger}=\eta^{-1}$ is unitary, which is consistent with providing a degenerate system of constraints that determine a contour in the complex plane (see Section 2).
Thus, we formulate the $i\phi^{3}$ theory in Eq. (3.1) in terms of a complex scalar field $\phi$, which has a ‘dual’ field $\tilde{\phi}^{\dagger}$. However, a complex scalar field has two propagating degrees of freedom rather than the one degree of freedom described by the original real scalar field. In the following subsection, we argue that the field $\phi$, while being complex, maintains only one degree of freedom. We do this by replacing ‘real’ numbers with ‘pseudo-real’ numbers.
3.3 Pseudo-reality
We propose that, for a pseudo-Hermitian Hamiltonian, ‘real’ fields are not equal to their Hermitian conjugate, i.e., $\phi^{\dagger}\neq\phi$. Instead, they are equal to their ‘dual’ field (3.16), making them ‘pseudo-real’:
$\tilde{\phi}^{\dagger}(x)=\phi(x)\>.$ (3.20)
This constraint is consistent with the Hamiltonian being $\hat{\eta}$-pseudo-Hermitian, i.e., $\hat{H}^{\dagger}=\hat{\eta}\hat{H}\hat{\eta}^{-1}$, and the field $\phi$, while being complex, has the the same number of degrees of freedom as the original real scalar field.
The pseudo-reality condition gives a non-trivial relationship between $\phi$ and its Hermitian conjugate $\phi^{\dagger}$:
$\hat{\phi}^{\dagger}(x)=\pi\hat{\eta}\hat{\phi}(x_{\eta})\hat{\eta}^{-1}=\pi\eta[\hat{\phi}(x)]\>,$ (3.21)
where we have used Eq. (3.19) relating the map $\eta$ to the operator $\hat{\eta}$.
Equation (3.21) presents a consistent resolution of the so-called Hermiticity Puzzle. Previous resolutions of this problem [14] have required a redefinition of Hermitian conjugation, where real scalar fields, such as $\phi$ here, are odd under Hermitian conjugation, i.e., anti-Hermitian. In the present context, we see that this redefinition of Hermitian conjugation amounts to associating it with the composition of Hermitian conjugation and $\eta$-conjugation.
The definition (3.21) of the pseudo-reality of the field $\phi$ is an operator version of the definition (2.4) of the pseudo-reality of a complex number $z$.
3.4 Parity and time-reversal
Let us assume that the field $\phi$ in the non-Hermitian Lagrangian (3.1) is complex and pseudo-real. This Lagrangian still has an anti-linear symmetry and is pseudo-Hermitian, but not with respect to the ‘standard’ parity and time-reversal transformations in Eqs. (3.3a) and (3.3b). Instead, we have a linear ‘parity-like’ and an anti-linear ‘time-reversal-like’ transformation, which map the field $\phi$ to its ‘dual’ conjugate $\tilde{\phi}$. For the ‘parity-like’ transformation $\eta_{P}$, we have:
$\eta_{P}:\>$ $\>\>\phi(x)\longmapsto\phi^{\eta_{P}}(x_{P})=\eta_{P}[\phi(x)]=-\tilde{\phi}(x)\>,\>\>\>\>\>\qquad\qquad\hat{\eta}_{P}\hat{\phi}(x_{P})\hat{\eta}_{P}^{-1}=-\hat{\tilde{\phi}}(x)\>,$ (3.22a)
$\mathscr{L}(x)\longmapsto\mathscr{L}^{\eta_{P}}(x_{P})=\mathscr{L}^{\dagger}(x)\>,\qquad\qquad\qquad\qquad\>\>\>\>\hat{\eta}_{P}\hat{\mathscr{L}}(x_{P})\hat{\eta}_{P}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)\>,$ (3.22b)
where the constant $\pi_{P}=-1$ in Eq. (3.16) reflects the pseudo-scalar nature of the field. For the ‘time-reversal-like’ transformation $\eta_{T}$, we have:
$\eta_{T}:\>$ $\>\>\phi(x)\longmapsto\phi^{\eta_{T}}(x_{T})=\eta_{T}[\phi(x)]=\tilde{\phi}(x)\>,\qquad\qquad\>\>\>\>\>\>\>\>\hat{\eta}_{T}\hat{\phi}(x_{T})\hat{\eta}_{T}^{-1}=\hat{\tilde{\phi}}(x)\>,$ (3.23a)
$\mathscr{L}(x)\longmapsto\mathscr{L}^{\eta_{T}}(x_{T})=\mathscr{L}^{\dagger}(x)\>,\qquad\qquad\qquad\qquad\>\>\>\>\hat{\eta}_{T}\hat{\mathscr{L}}(x_{T})\hat{\eta}_{T}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)\>,$ (3.23b)
where the constant is $\pi_{T}=+1$, as the anti-linearity of the time-reversal operator changes the sign of the complex interaction term in the Lagrangian. Hence, the Lagrangian is $\eta_{P}^{-1}\eta_{T}$-symmetric:
$\eta_{P}^{-1}\eta_{T}:\>\mathscr{L}(x)\longmapsto\mathscr{L}^{\eta_{P}^{-1}\eta_{T}}(x_{PT})=\mathscr{L}(x)\>,\>\>\>(\hat{\eta}_{P}^{-1}\hat{\eta}_{T})\hat{\mathscr{L}}(x_{PT})(\hat{\eta}_{P}^{-1}\hat{\eta}_{T})^{-1}=\hat{\mathscr{L}}(x)\>.$ (3.24)
However, in subsection $3.6$, we show that the operators $\hat{\eta}_{P}$ and $\hat{\eta}_{T}$ are related to the ‘standard’ parity $\hat{\mathcal{P}}$ and time-reversal $\hat{\mathcal{T}}$ transformations, such that $\hat{\eta}_{P}^{-1}\hat{\eta}_{T}=\hat{\mathcal{P}}\hat{\mathcal{T}}$, meaning the Lagrangian is still $PT$-symmetric with respect to the ‘standard’ $PT$ transformation.
The pseudo-reality condition in Eq. (3.20) relates the field $\phi$ to its Hermitian conjugate $\phi^{\dagger}$ via the linear operator $\hat{\eta}_{P}$:
$\hat{\phi}^{\dagger}(x)=-\hat{\eta}_{P}\hat{\phi}(x_{P})\hat{\eta}_{P}^{-1}=-\eta_{P}[\hat{\phi}(x)]\>,$ (3.25)
and via the anti-linear operator $\hat{\eta}_{T}$:
$\hat{\phi}^{\dagger}(x)=\hat{\eta}_{T}\hat{\phi}(x_{T})\hat{\eta}_{T}^{-1}=\eta_{T}[\hat{\phi}(x)]\>.$ (3.26)
As we will show in the following subsections, both of these conditions are equivalent and give a single constraint on the field $\phi$.
3.5 Finding $\eta$
In this subsection, we find the explicit form of the transformations $\eta_{P}$ and $\eta_{T}$. We begin with the ‘parity-like’ transformation $\eta_{P}$:
$\eta_{P}:\>\>\phi(x)\longmapsto\eta_{P}[\phi(x)]=-\tilde{\phi}(x)\>,$ (3.27)
where we have used the ‘pseudo-reality’ condition in Eq. (3.20). We see that the field $\eta_{P}[\phi(x)]$ solves the same equation of motion as the field $-\tilde{\phi}(x)$:
$\Box_{x}\eta_{P}[\phi(x)]+m^{2}\eta_{P}[\phi(x)]+\frac{ig}{2}\quantity(\eta_{P}[\phi(x)])^{2}=0\>.$ (3.28)
This equation is non-linear, but it can be solved perturbatively in orders of the coupling constant $g$. Thus, at order $g^{2}$, the map $\eta_{P}$ is of the form
$\eta_{P}[\phi(x)]=-\phi(x)+ig\int{\rm d}^{4}y\>G_{0}(x,y)\phi^{2}(y)+g^{2}\int{\rm d}^{4}y\>{\rm d}^{4}z\>G_{0}(x,y)\phi(y)G_{0}(y,z)\phi^{2}(z)+\mathcal{O}(g^{3})\>.$ (3.29)
Here, the two-point function $G_{0}(x,y)$ is the Green’s function (3.11) of the free Klein-Gordon equation. Note, however, that the field $\phi$ is not the free Klein-Gordon field $\phi_{0}$, but the complex interacting field of our Lagrangian, which itself can be expanded in orders of $g$ [see Eq. (3.10)]. The map $\eta_{P}$ bears some resemblance to the map derived in Refs. [22, 23]. [^1] In Ref. [22], it is suggested that there exists an arbitrariness to the Hermitian theory that corresponds to the $i\phi^{3}$ theory in its $PT$-symmetric regime. We disagree with this conclusion, and it is readily confirmed that the apparent arbitrariness remaining in Eq. (17) of Ref. [22] is eliminated when performing the path integral. This is to say that the terms proportional to the unconstrained coefficients $c_{2}$, $c_{4}$ and $c_{5}$ that appear therein cancel at order $g^{2}$.
The pseudo-reality condition (3.27) gives the relationship between the field $\phi$ and its Hermitian conjugate $\phi^{\dagger}$:
$\phi^{{\dagger}}(x)=\phi(x)-ig\int{\rm d}^{4}y\>G_{0}(x,y)\phi^{2}(y)-g^{2}\int{\rm d}^{4}y\>{\rm d}^{4}z\>G_{0}(x,y)\phi(y)G_{0}(y,z)\phi^{2}(z)+\mathcal{O}(g^{3})$ (3.30)
This equation defines a hypersurface in the complex field space and reduces the number of degrees of freedom of the complex field $\phi$ to that of a real field.
The time-reversal map $\eta_{T}$ solves the equation of motion for $-\eta_{P}$. Hence, the solution is straightforward
$\eta_{T}[\phi(x)]=\phi(x)-ig\int{\rm d}^{4}y\>G_{0}(x,y)\phi^{2}(y)-g^{2}\int{\rm d}^{4}y\>{\rm d}^{4}z\>G_{0}(x,y)\phi(y)G_{0}(y,z)\phi^{2}(z)\>+\mathcal{O}(g^{2})\;,$ (3.31)
and this yields the same pseudo-reality condition as Eq. (3.30).
3.6 $\eta$ operator
Now that we know how the transformations $\eta_{P}$ and $\eta_{T}$ act on the field $\phi$, we wish to find the corresponding Hermitian operators $\hat{\eta}_{P}$ and $\hat{\eta}_{T}$, with
$\eta_{P}[\hat{\phi}(x)]=\hat{\eta}_{P}\hat{\phi}(x_{P})\hat{\eta}^{-1}_{P}\qquad\text{and}\qquad\eta_{T}[\hat{\phi}(x)]=\hat{\eta}_{T}\hat{\phi}(x_{T})\hat{\eta}^{-1}_{T}\>.$ (3.32)
First, we need to choose a Green’s function $G_{0}(x,y)$ in the expansion (3.29). The equal-time commutation relations
$[\hat{\phi}(t,\vec{x}),\hat{\pi}(t,\vec{y})]=i\delta^{(3)}(\vec{x}-\vec{y})\>,$ (3.33a)
$[\hat{\phi}(t,\vec{x}),\hat{\phi}(t,\vec{y})]=[\hat{\pi}(t,\vec{x}),\hat{\pi}(t,\vec{y})]=0\>,$ (3.33b)
are only preserved for a Hermitian Green’s function. Here, the operator $\pi(x)$ is the conjugate momentum $\pi(x)=\frac{\partial\mathscr{L}(x)}{\partial\dot{\phi}(x)}=\dot{\phi}(x)$ of the interacting field $\phi$. However, we note that the field operator $\hat{\phi}(x)$ and its Hermitian conjugate $\hat{\phi}^{\dagger}(x)$ do not commute at equal times:
$[\hat{\phi}(t,\vec{x}),\hat{\phi}^{\dagger}(t,\vec{y})]\neq 0\>.$ (3.34)
It should not be surprising as $\phi^{\dagger}$ is not the dual field to $\phi$. Indeed, the dual field $\tilde{\phi}^{\dagger}$ commutes with $\phi$ at equal times:
$[\hat{\phi}(t,\vec{x}),\hat{\tilde{\phi}}^{{\dagger}}(t,\vec{y})]=0\>.$ (3.35)
The only Hermitian Green’s functions solving (3.11) are the retarded and the advanced propagators:
[…]
where $D(x,y)$ is the free Wightman function
$D(x,y)=\int\!\frac{{\rm d}^{3}\vec{p}}{(2\pi)^{3}}\>\frac{1}{2iE_{\vec{p}}}\,e^{-ip\cdot(x-y)}\>.$ (3.37)
In the remainder of this section, we choose to work with the retarded propagator. However, we could likewise use the advanced propagator.
We can show that the operator $\hat{\eta}_{P}=\hat{\eta}_{g}\hat{\mathcal{P}}$ can be written as a combination of the operator $\hat{\eta}_{g}$, yielding the series expansion in $g$, and the ‘standard’ parity operator $\hat{\mathcal{P}}$. The ‘standard’ parity operator has the usual form
$\hat{\mathcal{P}}(t)\hat{\phi}(t,\vec{x})\hat{\mathcal{P}}^{-1}(t)=-\hat{\phi}(t,-\vec{x})\>,\qquad\hat{\mathcal{P}}(t)=e^{\frac{\pi}{2}\int d^{3}z(\hat{\phi}(t,\vec{z})\hat{\pi}(t,-\vec{z})+\hat{\phi}(t,\vec{z})\hat{\pi}(t,\vec{z}))}\>.$ (3.38)
We note that the Lagrangian is pseudo-Hermitian with respect to $\hat{\eta}_{P}$, but not with respect to the $\hat{\mathcal{P}}$ operator.
The operator $\hat{\eta}_{g}$ is less trivial, and the full calculation is provided in Appendix A. We are looking for an expansion of the form
$\hat{\eta}_{g}(t,t_{0})\hat{\phi}(t,\vec{x})\hat{\eta}^{-1}_{g}(t,t_{0})$ $=\hat{\phi}(t,\vec{x})-ig\int^{t}_{t_{0}}{\rm d}y^{0}\int{\rm d}^{3}\vec{y}\>G_{R}(x,y)\hat{\phi}^{2}(y)$
$-g^{2}\int^{t}_{t_{0}}{\rm d}y^{0}\int^{y^{0}}_{t_{0}}{\rm d}z^{0}\int{\rm d}^{3}\vec{y}\>{\rm d}^{3}\vec{z}\>G_{R}(x,y)\hat{\phi}(y)G_{R}(y,z)\hat{\phi}^{2}(z)$
$+ig^{2}\int^{t}_{t_{0}}{\rm d}y^{0}\int^{y^{0}}_{t_{0}}{\rm d}z^{0}\int{\rm d}^{3}\vec{y}\>{\rm d}^{3}\vec{z}\>G_{R}(x,y)G_{R}(y,z)G_{R}(y,z)\hat{\phi}(z)\>.$ (3.39)
The last term is a quantum loop correction of order $\hbar$. As we take the classical limit $\hbar\mapsto 0$, this term vanishes and we recover the classical expansion in Eq. (3.29).
Given that the solution is similar to that of the interacting Klein-Gordon equation, we find that the series expansion is generated by an operator similar to the time-evolution operator; specifically,
$\hat{\eta}_{g}(t,t_{0})=\hat{U}^{-1}\quantity(-\frac{1}{3},t,t_{0})\>,$ (3.40)
where we define the operator $\hat{U}$ via
$\hat{U}\quantity(\alpha,t,t_{0})\coloneqq T\quantity[e^{\alpha g\int^{t}_{t_{0}}{\rm d}y^{0}\int{\rm d}^{3}\vec{y}\>\hat{\phi}^{3}(y^{0},\vec{y})}]\text{\>\>\>and\>\>\>}\hat{U}^{-1}\quantity(\alpha,t,t_{0})=\bar{T}\quantity[e^{-\alpha g\int^{t}_{t_{0}}{\rm d}y^{0}\int{\rm d}^{3}\vec{y}\>\hat{\phi}^{3}(y^{0},\vec{y})}]\>,$ (3.41)
for some constant $\alpha\in\mathbb{C}$. Here, $T$ is the time-ordering operator and $\bar{T}$ is the anti-time ordering operator. Note that $\hat{U}(\frac{1}{3!},t,t_{0})$ is (when replacing $\hat{\phi}$ by $\hat{\phi}_{0}$) the usual time-evolution operator whose action on the free Klein-Gordon field $\phi_{0}$ gives the series expansion in Eq. (3.10).
In Appendix B, we show that, at order $g^{2}$, the operators $\hat{\eta}_{g}$ and $\hat{\eta}_{P}$ are Hermitian:
[…]
and the ‘standard’ parity operator is $\hat{\eta}_{P}$-pseudo-Hermitian, i.e.,
$\hat{\mathcal{P}}^{\dagger}(t)=\hat{\eta}_{P}(t,t_{0})\hat{\mathcal{P}}(t)\hat{\eta}_{P}^{-1}(t,t_{0})\;.$ (3.43)
Hence, the ‘parity-like’ operator that makes our Lagrangian pseudo-Hermitian is given by
$\hat{\eta}_{P}(t,t_{0})=\hat{\eta}_{g}(t,t_{0})\>\hat{\mathcal{P}}(t)\>.$ (3.44)
In a similar manner, we can write the ‘time-reversal-like’ operator $\hat{\eta}_{T}=\hat{\eta}_{g}\hat{\mathcal{T}}$ as a combination of the ‘standard’ time-reversal operator
$\hat{\mathcal{T}}(t)\hat{\phi}(t,\vec{x})\hat{\mathcal{T}}^{-1}(t)=\hat{\phi}(-t,\vec{x})$ (3.45)
and the operator $\hat{\eta}_{g}$ in Eq. (3.40). Thus, the Lagrangian is anti-pseudo-Hermitian with respect to the anti-linear operator $\hat{\eta}_{T}$:
$\hat{\eta}_{T}(t,t_{0})=\hat{\eta}_{g}(t,t_{0})\hat{\mathcal{T}}(t).$ (3.46)
Given the Lagrangian is $\eta_{P}^{-1}\eta_{T}$-symmetric (3.24), we see that it is also $PT$-symmetric with respect to the ‘standard’ $PT$ transformation as $\hat{\eta}_{P}^{-1}\hat{\eta}_{T}=\hat{\mathcal{P}}\hat{\mathcal{T}}$.
Thus, we have demonstrated that by adopting a first-principles approach and introducing the pseudo-reality condition, we achieve a self-consistent formulation of the $i\phi^{3}$ theory. This framework resolves the contradictions arising from Euler-Lagrange equations and ‘standard’ parity and time-reversal transformations and ensures that the theory remains well-defined with a real spectrum under exact $PT$ symmetry.
4 Generalisation to higher spins
In the preceding sections, we have focussed on the simplest, spin-zero representation. We now turn our attention to higher-spin representations. Of particular interest is the coupling to gauge fields, and we begin from a well-studied theory composed of two complex scalar fields with an anti-Hermitian mass mixing [15, 17, 16]. Doing so, will provide an illustration of how pseudo-reality allows us to gauge non-Hermitian quantum field theories and to couple them to gravity consistently. Moreover, this two complex scalar field theory falls under the second approach to constructing non-Hermitian quantum field theories, constructed ab-initio from non-Hermitian composite operators. We will see, however, that pseudo-reality also allows to interpret this theory as the analytic continuation of a Hermitian theory.
4.1 Two complex scalar field theory
We start with the following free field theory (see, e.g., Refs. [15, 17, 16]):
$\mathcal{L}=\partial_{\mu}\tilde{\Phi}^{{\dagger}}\partial^{\mu}\Phi-\tilde{\Phi}^{{\dagger}}M^{2}\Phi\;,$ (4.1)
where the field $\Phi\equiv(\phi_{1},\phi_{2})$ is composed of two complex scalar fields $\phi_{1}$ and $\phi_{2}$. As the original Lagrangian, composed of $\Phi$ and $\Phi^{\dagger}$ is $PT$-symmetric and parity-pseudo-Hermitian, the dual field $\tilde{\Phi}^{{\dagger}}\equiv(\tilde{\phi}_{1}^{{\dagger}},\tilde{\phi}_{2}^{{\dagger}})$ is given via the parity transformation $\hat{\mathcal{P}}$
$\hat{\tilde{\Phi}}^{{\dagger}}(x)=\hat{\mathcal{P}}^{-1}\hat{\Phi}^{{\dagger}}(x_{P})\hat{\mathcal{P}}\>\Pi\;,\qquad\Pi=\begin{pmatrix}+1&0\\ 0&-1\end{pmatrix}\;.$ (4.2)
Here, $\Pi$ is the parity matrix, reflecting the scalar and pseudo-scalar nature of the fields $\phi_{1}$ and $\phi_{2}$ respectively. The mass matrix is skew-symmetric and $\Pi$-pseudo-Hermitian
$M^{2}=\begin{pmatrix}m_{1}^{2}&m_{5}^{2}\\ -m_{5}^{2}&m_{2}^{2}\end{pmatrix}\;,\qquad M^{2^{\dagger}}=\Pi M^{2}\Pi^{-1}\>,$ (4.3)
with $m_{1}^{2}$, $m_{2}^{2}$ and $m_{5}^{2}$ real and positive.
We will now show that the field $\hat{\Phi}$ is a complex field that can be decomposed into two pseudo-real fields. We start by writing
$\hat{\phi}_{j}=\frac{1}{\sqrt{2}}\Big{(}\hat{\chi}_{j1}+i\hat{\chi}_{j2}\Big{)}\;,\qquad\text{and}\qquad\hat{\tilde{\phi}}^{{\dagger}}_{j}=\frac{1}{\sqrt{2}}\Big{(}\hat{\tilde{\chi}}^{{\dagger}}_{j1}-i\hat{\tilde{\chi}}^{{\dagger}}_{j2}\Big{)}\;.$ (4.4)
Here, we take the fields $\hat{\chi}_{jk}\in\mathbb{C}$, following an approach similar to Ref. [14] (cf. the alternative approach of Ref. [24], where the field is decomposed into two real components). If the field components $\hat{\chi}_{jk}$ and $\hat{\tilde{\chi}}_{jk}^{{\dagger}}$ are unrelated to each other, then we have eight degrees of freedom, rather than the four that we would expect for two complex scalar fields. We recover four degrees of freedom by imposing pseudo-reality on the field components $\hat{\chi}_{jk}$, namely that
$\hat{\tilde{\chi}}^{{\dagger}}_{jk}(x)=\hat{\chi}_{jk}(x)\;.$ (4.5)
The Lagrangian then takes the form
^$\hat{\mathcal{L}}=\frac{1}{2}\sum_{j,k}\partial_{\mu}\hat{\chi}_{jk}\partial^{\mu}\hat{\chi}_{jk}-\frac{1}{2}m_{1}^{2}\left(\hat{\chi}_{11}^{2}+\hat{\chi}_{12}^{2}\right)-\frac{1}{2}m_{2}^{2}\left(\hat{\chi}_{22}^{2}+\hat{\chi}_{21}^{2}\right)-im_{5}^{2}\left(\hat{\chi}_{11}\hat{\chi}_{22}-\hat{\chi}_{12}\hat{\chi}_{21}\right)\;,$ (4.6)
and the resulting equations of motion for the fields $\hat{\chi}_{jk}$ are
$\Box\hat{\chi}_{11}+m_{1}^{2}\hat{\chi}_{11}+im_{5}^{2}\hat{\chi}_{22}=0\;,$ (4.7a)
$\Box\hat{\chi}_{12}+m_{1}^{2}\hat{\chi}_{12}-im_{5}^{2}\hat{\chi}_{21}=0\;,$ (4.7b)
$\Box\hat{\chi}_{21}+m_{2}^{2}\hat{\chi}_{21}-im_{5}^{2}\hat{\chi}_{12}=0\;,$ (4.7c)
$\Box\hat{\chi}_{22}+m_{2}^{2}\hat{\chi}_{22}+im_{5}^{2}\hat{\chi}_{11}=0\;.$ (4.7d)
This theory is analytically continued from a Hermitian theory by taking $m_{5}^{2}\to im_{5}^{2}$, and we see immediately, as for the $i\phi^{3}$ theory discussed in Sec. 3, that the fields $\hat{\chi}_{jk}$ cannot be real. We now need only to ensure the self-consistency of the pseudo-reality of the fields $\hat{\chi}_{jk}$.
The pseudo-reality conditions for the components $\hat{\chi}_{jk}$ are given via the parity transformation:
$\hat{\tilde{\chi}}^{{\dagger}}_{11}$ $=\hat{\mathcal{P}}^{-1}\hat{\chi}^{\dagger}_{11}\hat{\mathcal{P}}=\hat{\chi}_{11}\;,$ (4.8a)
$\hat{\tilde{\chi}}^{{\dagger}}_{12}$ $=\hat{\mathcal{P}}^{-1}\hat{\chi}^{\dagger}_{12}\hat{\mathcal{P}}=\hat{\chi}_{12}\;,$ (4.8b)
$\hat{\tilde{\chi}}^{{\dagger}}_{21}$ $=-\hat{\mathcal{P}}^{-1}\hat{\chi}^{\dagger}_{21}\hat{\mathcal{P}}=\hat{\chi}_{21}\;,$ (4.8c)
$\hat{\tilde{\chi}}^{{\dagger}}_{22}$ $=-\hat{\mathcal{P}}^{-1}\hat{\chi}^{\dagger}_{22}\hat{\mathcal{P}}=\hat{\chi}_{22}\;.$ (4.8d)
Note that, unlike the case of the $i\phi^{3}$ theory, where the degree of non-Hermiticity is controlled by the small coupling constant $g$, we cannot obtain an expression $\eta[\hat{\chi}_{jk}]$ in terms of the original components $\hat{\chi}_{jk}$.
4.2 Gauged model
A number of attempts have been made in the literature to gauge similar scalar models that incorporate two complex scalar fields [25, 26]. There are two possibilities: (i) couple the gauge field to a Hermitian but non-conserved current or (ii) couple the gauge field to a conserved but non-Hermitian current. Method (i) leads to an inconsistent Maxwell equation and breaks the gauge invariance of the Lagrangian. Method (ii) is in contradiction with the reality of the gauge field. We will show in this section that method (ii) is consistent so long as the gauge field is taken to be a pseudo-real vector field.
To see this, we minimally couple a U(1) gauge field to the doublet $\Phi=(\phi_{1},\phi_{2})$ of two complex scalar fields as follows:
$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}\tilde{F}^{{\dagger}\mu\nu}-\frac{1}{2\xi}(\partial_{\mu}A^{\mu})(\partial_{\mu}\tilde{A}^{{\dagger}\mu})+\tilde{D}_{\mu}^{{\dagger}}\tilde{\Phi}^{{\dagger}}D^{\mu}\Phi-\tilde{\Phi}^{{\dagger}}M^{2}\Phi\;,$ (4.9)
where $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the usual field strength tensor, and the gauge covariant derivatives are
$D_{\mu}=\partial_{\mu}-ieA_{\mu}\;,\qquad\tilde{D}_{\mu}^{{\dagger}}=\partial_{\mu}+ie\tilde{A}_{\mu}^{{\dagger}}\;,$ (4.10)
Consider the operator-valued conserved current
$\hat{j}^{\nu}=ie\left[\hat{\tilde{\Phi}}^{{\dagger}}\overset{\rightarrow}{D^{\nu}}\hat{\Phi}-\hat{\tilde{\Phi}}^{{\dagger}}\overset{\leftarrow}{\tilde{D}^{\nu{\dagger}}}\hat{\Phi}\right]\;.$ (4.11)
This current can be written in terms of the pseudo-real components as follows:
$\hat{j}^{\nu}$ $=-e\left[\hat{\chi}_{11}\overset{\rightarrow}{\partial^{\nu}}\hat{\chi}_{12}-\hat{\chi}_{11}\overset{\leftarrow}{\partial^{\nu}}\hat{\chi}_{12}+\hat{\chi}_{21}\overset{\rightarrow}{\partial^{\nu}}\hat{\chi}_{22}-\hat{\chi}_{21}\overset{\leftarrow}{\partial^{\nu}}\hat{\chi}_{22}\right]$
$\phantom{=}+\frac{1}{2}e^{2}\left(\hat{A}^{\nu}+\hat{\tilde{A}}^{\nu{\dagger}}\right)\left[\hat{\chi}_{11}^{2}+\hat{\chi}_{12}^{2}+\hat{\chi}_{21}^{2}+\hat{\chi}_{22}^{2}\right]\;.$ (4.12)
This current is pseudo-real if the gauge field satisfies the pseudo-reality condition
$\hat{\tilde{A}}^{\mu{\dagger}}(x)=\hat{\mathcal{P}}^{-1}\hat{A}^{\nu{\dagger}}(x_{P})\hat{\mathcal{P}}\Pi_{\nu}^{\phantom{\nu}\mu}=\hat{A}^{\mu}(x)\;,$ (4.13)
where $\Pi_{\nu}^{\phantom{\nu}\mu}=\delta_{\nu}^{\phantom{\nu}\mu}$ in this case.
The following Maxwell equations are now all internally consistent:
$\Box\hat{A}^{\mu}-\left(1-\frac{1}{\xi}\right)\partial^{\nu}\partial_{\mu}\hat{A}^{\mu}=\hat{j}^{\mu}\;,$ (4.14a)
$\Box\hat{\tilde{A}}^{\mu}-\left(1-\frac{1}{\xi}\right)\partial^{\nu}\partial_{\mu}\hat{\tilde{A}}^{\mu}=\hat{\tilde{j}}^{\mu}\;,$ (4.14b)
$\Box\hat{A}^{\mu{\dagger}}-\left(1-\frac{1}{\xi}\right)\partial^{\nu}\partial_{\mu}\hat{A}^{\mu{\dagger}}=\hat{j}^{\mu{\dagger}}\,,$ (4.14c)
and the pseudo-reality condition (4.13) implies that the photon field $A_{\mu}$, although complex, possesses only two propagating degrees of freedom.
We anticipate that this consistent pseudo-Hermitian formulation of a U(1) gauge theory will allow to revisit existing discrepancies in the treatments of the Higgs mechanism in similar non-Hermitian field theories (cf. Refs. [14, 25, 27]). This conjecture may be checked in future work.
4.3 Coupling to gravity
Having successfully gauged the two complex scalar model in the previous section, it is now clear how to couple non-Hermitian field theories to spin-2 fields. Consider, e.g., the free scalar sector described above. We can couple this theory minimally to Einstein gravity through the matter action
$S=\int{\rm d}^{4}x\sqrt{-g}\left[g^{\mu\nu}\partial_{\mu}\tilde{\Phi}^{{\dagger}}\partial_{\nu}\Phi-\tilde{\Phi}^{{\dagger}}M^{2}\Phi\right]\;,$ (4.15)
where $g_{\mu\nu}$ is the metric tensor. The resulting energy-momentum tensor is
$T^{\mu\nu}=\frac{1}{2}\partial^{(\mu}\tilde{\Phi}^{{\dagger}}\partial^{\nu)}\Phi-g^{\mu\nu}\mathcal{L}\;,$ (4.16)
where the parenthesis indicates symmetrisation of the Lorentz indices. This energy-momentum tensor is not Hermitian, but it is pseudo-Hermitian with respect to the transformation $\eta$, with $\eta=\mathcal{P}$ in this case. The Einstein equation
$G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=\kappa T_{\mu\nu}$ (4.17)
then implies that the Einstein tensor $G_{\mu\nu}$ is not Hermitian in general. Nevertheless, the Einstein equation is consistent as long as the metric $g_{\mu\nu}$, Ricci scalar $R$ and Ricci tensor $R_{\mu\nu}$ are pseudo-real with respect to the same transformation $\eta$, i.e.,
$\tilde{R}^{{\dagger}}_{\mu\nu}-\frac{1}{2}\tilde{g}_{\mu\nu}^{{\dagger}}\tilde{R}^{{\dagger}}=R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R\;,$ (4.18)
with the metric satisfying
$\tilde{g}^{{\dagger}}_{\mu\nu}(x)=g_{\mu\nu}(x)\;,$ (4.19)
where, at the operator level, we would have $\hat{\tilde{g}}^{{\dagger}}_{\mu\nu}(x)=\hat{\eta}^{-1}\hat{g}^{{\dagger}}_{\alpha\beta}(x_{\eta})\hat{\eta}\Pi^{\alpha\beta}_{\mu\nu}$, with $\Pi^{\alpha\beta}_{\mu\nu}$ fixed by the pseudo-Hermiticity of the Lagrangian.
Most notably, the pseudo-reality of the curvature tensors (4.18) and the metric (4.19) is determined by the pseudo-Hermiticity properties of all non-Hermitian sectors to which they couple, as was the case for the vector field in the previous section. We leave further study of these pseudo-real geometries to future work, but we anticipate that the observations in this work pave the way to consistent formulations of non-Hermitian cosmological models.
5 Conclusions
We have presented a general framework within which Hermiticity puzzles in pseudo-Hermitian quantum field theories can be resolved. We refer to this property as pseudo-reality. We first introduced the concept of pseudo-reality for complex numbers and then generalized it to quantum fields, promoting a real field defined on the real axis to a pseudo-real field defined on a contour in the complex plane. This approach resolves a long-standing inconsistency in the minimal coupling of pseudo-Hermitian sectors to gauge fields. We have focused here on a U(1) gauge theory and indicated how pseudo-reality enables a consistent coupling to gravity. Interestingly, the pseudo-reality conditions depend on the map $\eta$ that determines the pseudo-Hermiticity of the full Hamiltonian. Thus, the pseudo-reality condition of a given bosonic field is determined by the non-Hermitian sectors to which it couples, making clear the subtleties of treating interacting pseudo-Hermitian quantum field theories. We leave in-depth study of such theories and their couplings to gravity to future work.
Acknowledgements.
[…]
Data Access Statement
No data were created or analysed in this study.
Appendix A Calculating the $\hat{\eta}$ operator
We note that the expansion of the $\eta_{P}$ transformation in Eq. (3.29) has the same terms as the expansion of $\phi$ in Eq. (3.10). Hence, we make an Ansatz that the operator $\hat{\eta}_{P}$ should be of the form of the time-evolution operator, but with differing coefficients and expressed in terms of the full, interacting field $\hat{\phi}$. We define an operator $\hat{U}$
$\hat{U}\quantity(\alpha,t,t_{0})\coloneqq T\quantity[e^{\alpha g\int^{t}_{t_{0}}{\rm d}y^{0}\int{\rm d}^{3}\vec{y}\>\hat{\phi}^{3}(y^{0},\vec{y})}]\text{\>\>\>and\>\>\>}\hat{U}^{-1}\quantity(\alpha,t,t_{0})=\bar{T}\quantity[e^{-\alpha g\int^{t}_{t_{0}}{\rm d}y^{0}\int{\rm d}^{3}\vec{y}\>\hat{\phi}^{3}(y^{0},\vec{y})}]\>,$ (A.1)
for some constant $\alpha\in\mathbb{C}$. Here, $T$ is the time-ordering operator and $\bar{T}$ is the anti-time ordering operator.
We proceed by expanding
$\hat{U}^{-1}(\alpha,t,t_{0})\hat{\phi}(t,\vec{x})\hat{U}(\alpha,t,t_{0})=\hat{\phi}_{x}-\alpha g\int_{t_{0},\>\vec{y}}^{x^{0}}\>\quantity[\hat{\phi}^{3}_{y},\hat{\phi}_{x}]+\alpha^{2}g^{2}\int_{t_{0},\>\vec{y}}^{x^{0}}\>\int_{t_{0},\>\vec{z}}^{y^{0}}\>\quantity[\hat{\phi}^{3}_{z},\quantity[\hat{\phi}^{3}_{y},\hat{\phi}_{x}]]\;,$ (A.2)
[…]
We note that the commutator $[\hat{\phi}_{x},\hat{\phi}_{y}]$ corresponds to that of the interacting field $\hat{\phi}$ and not to that of the free Klein-Gordon field $\hat{\phi}_{0}$. However, we can expand the interacting field as a series in the coupling constant $g$. For our purposes, we only need to perform this expansion at first order in $g$:
$\hat{\phi}_{x}=\hat{\phi}_{0x}+\frac{ig}{2}\int_{t_{0},\>\vec{y}}^{x^{0}}\>G_{Rxy}\hat{\phi}_{0y}^{2}+\mathcal{O}(g^{2})\;.$ (A.3)
This relation can be inverted (again, at linear order in $g$) to obtain
$\hat{\phi}_{0x}=\hat{\phi}_{x}-\frac{ig}{2}\int_{t_{0},\>\vec{y}}^{x^{0}}\>G_{Rxy}\hat{\phi}_{y}^{2}+\mathcal{O}(g^{2})\;.$ (A.4)
Herein, $G_{Rxy}\coloneqq G_{R}(x,y)$ denotes the retarded propagator. We also denote the Pauli–Jordan function by $G_{xy}\coloneqq D_{xy}-D_{yx}$, which only becomes the retarded propagator with appropriate integral limits.
We calculate the commutators for the expansion in Eq. (A.2) at order $g$:
$[\hat{\phi}_{y},\hat{\phi}_{x}]$ $=\quantity[\hat{\phi}_{0y}+\frac{ig}{2}\int_{t_{0},\>\vec{z}}^{y^{0}}\>G_{Ryz}\hat{\phi}_{0z}^{2}+\mathcal{O}(g^{2}),\hat{\phi}_{0x}+\frac{ig}{2}\int_{t_{0},\>\vec{z}}^{x^{0}}\>G_{Rxz}\hat{\phi}_{0z}^{2}+\mathcal{O}(g^{2})]$
$=\quantity[\hat{\phi}_{0y},\hat{\phi}_{0x}]-g\int_{t_{0},\>\vec{z}}^{x^{0}}\>G_{Rxz}G_{yz}\hat{\phi}_{0z}+g\int_{t_{0},\>\vec{z}}^{y^{0}}\>G_{xz}G_{Ryz}\hat{\phi}_{0z}+\mathcal{O}(g^{2})\;,$ (A.5)
and invert the expression for $\phi$ to obtain
$[\hat{\phi}_{y},\hat{\phi}_{x}]=-iG_{xy}-g\int_{t_{0},\>\vec{z}}^{x^{0}}\>G_{Rxz}G_{yz}\hat{\phi}_{z}+g\int_{t_{0},\>\vec{z}}^{y^{0}}\>G_{xz}G_{Ryz}\hat{\phi}_{z}\;.$ (A.6)
Hereafter, we omit the terms that are subleading in the coupling constant $g$.
The last relation allows us to calculate the commutator:
$[\hat{\phi}^{3}_{y},\hat{\phi}_{x}]$ $=-3iG_{xy}\hat{\phi}^{2}_{y}-3g\int_{t_{0},\>\vec{z}}^{x^{0}}\>G_{Rxz}G_{yz}\hat{\phi}_{z}\hat{\phi}^{2}_{y}+3g\int_{t_{0},\>\vec{z}}^{y^{0}}\>G_{xz}G_{Ryz}\hat{\phi}_{z}\hat{\phi}^{2}_{y}$
$\phantom{=}-3ig\int_{t_{0},\>\vec{z}}^{x^{0}}\>G_{Rxz}G_{yz}G_{yz}\hat{\phi}_{y}+3ig\int_{t_{0},\>\vec{z}}^{y^{0}}\>G_{xz}G_{Ryz}G_{yz}\hat{\phi}_{y}\;.$ (A.7)
For the third term in Eq. (A.2), we only need the commutator at 00-th order in g:
$[\hat{\phi}^{3}_{z},[\hat{\phi}^{3}_{y},\hat{\phi}_{x}]]=-18G_{xy}G_{yz}\hat{\phi}_{y}\hat{\phi}^{2}_{z}+18iG_{xy}G_{yz}G_{yz}\hat{\phi}_{z}\;.$ (A.8)
Plugging the commutators into Eq. (A.2), we get
$\hat{U}^{-1}(\alpha,t,t_{0})\hat{\phi}(t,\vec{x})\hat{U}(\alpha,t,t_{0})$ $=\hat{\phi}_{x}+3i\alpha g\int_{t_{0},\>\vec{y}}^{x^{0}}\>G_{Rxy}\hat{\phi}^{2}_{y}$
$\phantom{=}-3g^{2}\alpha(1+6\alpha)\int_{t_{0},\>\vec{y}}^{x^{0}}\>\int_{t_{0},\>\vec{z}}^{y^{0}}\>G_{Rxy}G_{Ryz}\hat{\phi}_{y}\hat{\phi}^{2}_{z}$
$\phantom{=}+3i\alpha g^{2}(1+6\alpha)\int_{t_{0},\>\vec{y}}^{x^{0}}\>\int_{t_{0},\>\vec{z}}^{y^{0}}\>G_{xy}G_{yz}G_{yz}\hat{\phi}_{z}\;.$ (A.9)
Finally, taking $\alpha=-\frac{1}{3}$, we get the operator $\hat{\eta}_{g}(t)=\hat{U}^{-1}\quantity(-\frac{1}{3},t,t_{0})$ in Eq. (3.6)
$\hat{\eta}_{g}(t,t_{0})\hat{\phi}(t,\vec{x})\hat{\eta}_{g}^{-1}(t,t_{0})$ $=\hat{\phi}_{x}-ig\int_{t_{0},\>\vec{y}}^{x^{0}}\>G_{Rxy}\hat{\phi}^{2}_{y}$
$\phantom{=}-g^{2}\int_{t_{0},\>\vec{y}}^{x^{0}}\>\int_{t_{0},\>\vec{z}}^{y^{0}}\>G_{Rxy}G_{Ryz}\hat{\phi}_{y}\hat{\phi}^{2}_{z}$
$\phantom{=}+ig^{2}\int_{t_{0},\>\vec{y}}^{x^{0}}\>\int_{t_{0},\>\vec{z}}^{y^{0}}\>G_{xy}G_{yz}G_{yz}\hat{\phi}_{z}\;,$ (A.10)
with the last term appearing due to quantum loop corrections. It is proportional to $\hbar$ and vanishes in the classical limit $\hbar\mapsto 0$.
Appendix B Proving the Hermiticity of $\hat{\eta}$
In this Appendix, we show that $\hat{\eta}_{P}$ and $\hat{\eta}_{g}$ are Hermitian while $\hat{\mathcal{P}}$ is $\hat{\eta}_{P}$-pseudo-Hermitian.
We expand the operator $\hat{U}$ at order $g^{2}$:
$\hat{U}(\alpha,t,t_{0})=T\quantity[e^{\alpha\int_{t_{0},\>\vec{y}}^{x^{0}}\>\hat{\phi}^{3}_{y}}]=1+\alpha g\int_{t_{0},\>\vec{y}}^{x^{0}}\>\hat{\phi}^{3}_{y}+\alpha^{2}g^{2}\int_{t_{0},\>\vec{y}}^{x^{0}}\>\int_{t_{0},\>\vec{z}}^{y^{0}}\>\hat{\phi}^{3}_{y}\hat{\phi}^{3}_{z}\;.$ (B.1)
Taking its Hermitian conjugate, we have
$\hat{U}^{{\dagger}}(\alpha,t,t_{0})=1+\alpha g\int_{t_{0},\>\vec{y}}^{x^{0}}\>\hat{\phi}^{3^{\dagger}}_{y}+\alpha^{2}g^{2}\int_{t_{0},\>\vec{y}}^{x^{0}}\>\int_{t_{0},\>\vec{z}}^{y^{0}}\>\hat{\phi}^{3^{\dagger}}_{z}\hat{\phi}^{3^{\dagger}}_{y}\;.$ (B.2)
The pseudo-reality condition gives us an expansion of $\hat{\phi}^{\dagger}$ in terms of $\hat{\phi}$ [see Eq. (3.30)]. At order $g$, we have
$\hat{\phi}^{\dagger}_{x}=\hat{\phi}_{x}-ig\int_{t_{0},\>\vec{y}}^{x^{0}}\>G_{Rxy}\hat{\phi}^{2}_{y}\;.$ (B.3)
Hence, the interaction term, commuting all $\hat{\phi}_{x}$ to the left of $\hat{\phi}_{y}$, is
$\hat{\phi}^{3^{\dagger}}_{x}=\hat{\phi}^{3}_{x}-3ig\int_{t_{0},\>\vec{y}}^{x^{0}}\>G_{Rxy}\hat{\phi}^{2}_{x}\hat{\phi}^{2}_{y}-6g\int_{t_{0},\>\vec{y}}^{x^{0}}\>G_{Rxy}G_{xy}\hat{\phi}_{x}\hat{\phi}_{y}+2ig\int_{t_{0},\>\vec{y}}^{x^{0}}\>G_{Rxy}G_{xy}G_{xy}\;.$ (B.4)
We insert this expression into Eq. (B.2) and switch the order of $\hat{\phi}^{3}_{z}\hat{\phi}^{3}_{y}$ in the last term using the commutator
$[\hat{\phi}^{3}_{z},\hat{\phi}^{3}_{y}]=-9iG_{yz}\hat{\phi}^{2}_{y}\hat{\phi}^{2}_{z}-18G_{yz}G_{yz}\hat{\phi}_{y}\hat{\phi}_{z}+6iG_{yz}G_{yz}G_{yz}\;.$ (B.5)
Thus, we get $\hat{U}^{\dagger}$ at order $g^{2}$
$\hat{U}^{\dagger}(\alpha,t,t_{0})$ $=1+\alpha g\int_{t_{0},\>\vec{y}}^{x^{0}}\>\hat{\phi}^{3}_{y}+\alpha^{2}g^{2}\int_{t_{0},\>\vec{y}}^{x^{0}}\>\int_{t_{0},\>\vec{z}}^{y^{0}}\>\hat{\phi}^{3}_{y}\hat{\phi}^{3}_{z}$
$\phantom{=}-3i\alpha g^{2}(1+3\alpha)\int_{t_{0},\>\vec{y}}^{x^{0}}\>\int_{t_{0},\>\vec{z}}^{y^{0}}\>G_{Ryz}\hat{\phi}^{2}_{y}\hat{\phi}^{2}_{z}$
$\phantom{=}-6\alpha g^{2}(1+3\alpha)\int_{t_{0},\>\vec{y}}^{x^{0}}\>\int_{t_{0},\>\vec{z}}^{y^{0}}\>G_{Ryz}G_{yz}\hat{\phi}_{y}\hat{\phi}_{z}$
$\phantom{=}+2i\alpha g^{2}(1+3\alpha)\int_{t_{0},\>\vec{y}}^{x^{0}}\>\int_{t_{0},\>\vec{z}}^{y^{0}}\>G_{yz}G_{yz}G_{yz}\;.$ (B.6)
For the operator $\hat{\eta}_{g}$, we have $\alpha=-\frac{1}{3}$ and all the loop terms cancel leaving us with
$\hat{\eta}_{g}^{{-1}^{\dagger}}=1+\alpha g\int_{t_{0},\>\vec{y}}^{x^{0}}\>\hat{\phi}^{3}_{y}+\alpha^{2}g^{2}\int_{t_{0},\>\vec{y}}^{x^{0}}\>\int_{t_{0},\>\vec{z}}^{y^{0}}\>\hat{\phi}^{3}_{y}\hat{\phi}^{3}_{z}=T\quantity[e^{\alpha g\int_{t_{0},\>\vec{y}}^{x^{0}}\>\hat{\phi}^{3}_{y}}]=\hat{\eta}_{g}^{-1}\;.$ (B.7)
Hence, at order $g^{2}$, the operator $\hat{\eta}_{g}^{\dagger}=\hat{\eta}_{g}$ is Hermitian.
We can also show that the ‘standard’ parity operator
$\hat{\mathcal{P}}(t)=e^{\frac{\pi}{2}\int d^{3}z\>\hat{\phi}(t,\vec{z})\hat{\pi}(t,-\vec{z})+\hat{\phi}(t,\vec{z})\hat{\pi}(t,\vec{z})}$ (B.8)
is $\hat{\eta}_{P}$-pseudo-Hermitian. Taking the Hermitian conjugate, we have
$\hat{\mathcal{P}}^{\dagger}(t)=e^{\frac{\pi}{2}\int d^{3}z\>\hat{\phi}^{\dagger}(t,\vec{z})\hat{\pi}^{\dagger}(t,-\vec{z})+\hat{\phi}^{\dagger}(t,\vec{z})\hat{\pi}^{\dagger}(t,\vec{z})}\;,$ (B.9)
where we have used the fact that $\hat{\phi}$ and $\hat{\pi}$ commute at equal times up to a delta function, giving an over all phase to the operator $\hat{\mathcal{P}}$.
Using the expansion of $\hat{\phi}^{\dagger}$ in terms of the field $\hat{\phi}$, we can show that $\hat{\pi}^{\dagger}$ satisfies an analogous pseudo-reality condition:
$\hat{\phi}^{\dagger}(x)=-\hat{\eta}_{P}(t)\hat{\phi}(x_{P})\hat{\eta}_{P}^{-1}(t)\qquad\text{and}\qquad\hat{\pi}^{\dagger}(x)=-\hat{\eta}_{P}(t)\hat{\pi}(x_{p})\hat{\eta}_{P}(t)\;.$ (B.10)
Hence, using the pseudo-reality condition, we find that
$\hat{\mathcal{P}}^{\dagger}(t)=\hat{\eta}_{P}(t,t_{0})\hat{\mathcal{P}}(t)\hat{\eta}^{-1}_{P}(t,t_{0})\,,$ (B.11)
implying that the operator $\hat{\mathcal{P}}$ is $\hat{\eta}_{P}$-pseudo-Hermitian.
Finally, we can use the above results to show that $\hat{\eta}_{P}=\hat{\eta}_{g}\hat{\mathcal{P}}$ is Hermitian:
$\hat{\eta}_{P}^{\dagger}=\hat{\mathcal{P}}^{\dagger}\hat{\eta}_{g}^{\dagger}=\hat{\eta}_{P}\hat{\mathcal{P}}\hat{\eta}_{P}^{-1}\hat{\eta}_{g}=\hat{\eta}_{P}\hat{\mathcal{P}}\hat{\mathcal{P}}^{-1}\hat{\eta}_{g}^{-1}\hat{\eta}_{g}=\hat{\eta}_{P}\;.$ (B.12)